OpenStudy (anonymous):

Help please!!! If 2x + 5y = 21 and x + y = 3, then what is the value of x? It is your choice to use either substitution or elimination,

OpenStudy (anonymous):

and can you please explain it to me!!

OpenStudy (mathstudent55):

This is a system of equations. there are several methods for solving a system of equations. I will show you the method called elimination.

OpenStudy (mathstudent55):

In the elimination method, you always end up adding the equations in order to eliminate one valriable.

OpenStudy (mathstudent55):

If you look at your equations, you have 2x and x. Added together, that is 3x. that is not eliminated. Now if you look at the two y-terms you have 5y and y. Once again, when you add 5y and y, you get 6y, which means y is not eliminated.

OpenStudy (mathstudent55):

You need to do something to one equations or to both equations so that when you add the equations together, one variable will add to zero.

OpenStudy (mathstudent55):

The way you do that is to multiply one equation or both equations by a number or numbers. You choose the number or numbers carefully in a way that ensures one variable will be eliminated when the equations are added together.

OpenStudy (anonymous):

what is happening? is it a lag?

OpenStudy (aa1602):

2x+5y=21 2x+5y =21 2*( 3-y) +5y =21 6-2y+5y=21 => 6+3y=21 x+y=3 => x=3-y x=3-y x=3-y => x=3-y 3y=21-6 => 3y=15 => y= 15/3 => y=5 => y=5 x=3-y =>x=3-y => x=3-y => x=3-5 => x= -2 As you can see you just substitute at the top the x for 3-y, so you get an equation with one unknown variable, and when you get the value of Y, you put it at the other equation , to get the value of x. I hope this makes sense :)

OpenStudy (anonymous):

okay i get half of it but the question?

OpenStudy (aa1602):

Well, the value of x appears to be -2. If thats what you're looking for.

OpenStudy (mathstudent55):

2x + 5y = 21 x + y = 3 Since you are solving only for x, we need to eliminate y. If you multiply the second equation by -5, the y term of the second equation will become -5y. Then when you add that -5y term to the 5y of the first equation, you eliminate the y terms leaving only x. 2x + 5y = 21 --- (just copy the equation) ----> 2x + 5y = 21 x + y = 3 --- (multiply entire eq. by -5) ---> -5x - 5y = -15 (add) -------------- -3x = 6 Once you add the equations, you get -3x = 6. This equation can now be easily solved for x. Divide both sides by -3: x = -2

OpenStudy (anonymous):

ohhhh